OpenStudy (anonymous):

what is the solution to the proportion 12/15 = 8/x ?

Mehek (mehek14):

multiply 1/8 x 12/15

OpenStudy (welshfella):

cross multiply 12x = 15*8

Mehek (mehek14):

\(\dfrac{1}{8}{*}\dfrac{12}{15}\)

OpenStudy (anonymous):

so it would be ten??

Mehek (mehek14):

yes 12/15 = 0.8 8/10 = 0.8

OpenStudy (anonymous):

awsome tnxs

Mehek (mehek14):

yw